OpenStudy (anonymous):

The Boron has an atomic mass of 10.811 u. Calculate the percentage of natural abundance for both isotopes, 10B and 11B. The isotopic mass for 10B is 10.012937 and for 11B is 11.009305.

OpenStudy (anonymous):

You need to setup two formulas for this: Let 10B=x and let 11B=y Both isotopes add up to 100% \[x+y=1\] \[x=1-y\] We know that a combination of both based on percentage = 10.811 \[10.012937x+11.009305y=10.811\] Substitute: \[10.012937(1-y)+11.009305y=10.811\] \[10.012937-10.012937y+11.009305y=10.811\] Combine like types: \[0.996368y=0.798063\] \[y=\frac{0.798063}{0.996368}\] \[y=0.800972\] Based on this we conclude that 10B is 20% and 11B is 80%.

OpenStudy (anonymous):

Thank you, I solved it like 10 minutes later, but thanks for your answer.